Four tickets marked 00,01,10, 11 respectively are placed in a bag. A ticket is drawn at random five times, being replaced each time. Find the probability that the sum of the numbers on the tickets thus drawn is 23.


1
Expert's answer
2021-07-20T14:26:35-0400

The total number of ways to draw 4 tickets, five times "=4^5 ." This is the total number of possible events, that is, the number of sample spaces.

Now, possible outcomes for the sum of the number of tickets are "(11, 11, 01, 00, 00)," "(11, 10, 01, 01, 00), (10, 10, 01, 01, 01)."

Now, we have to find the number of ways to arrange these tickets because it is possible that they may occur in different order.

Therefore, there are a number of ways to arrange "(11, 11, 01, 00, 00)"


"\dfrac{5!}{2!\times2!},"

because both "11" and "00" are occurring twice.

Number of ways to arrange "(11, 10, 01, 01, 00)"is


"\dfrac{5!}{2!},"

because "00" is occurring twice.

Number of ways to arrange "(10, 10, 01, 01, 01)"is


"\dfrac{5!}{2!\times3!},"

because "10" is occurring twice and "01" is occurring thrice.

Therefore, total number of ways to arrange these outcomes


"\dfrac{5!}{2!\times2!}+\dfrac{5!}{2!}+\dfrac{5!}{2!\times3!}"

"=\dfrac{120(3+6+1)}{12}=100"

The probability that the sum of the number on tickets drawn is 23 will be

"\dfrac{100}{4^5}=\dfrac{25}{256}"